# : Accelerating the Science of Language Models

### v3

@@ Appendix B Adaptation Training Details
After instruction finetuning, we then use the following hyperparameters for DPO training, following Ivison et al. [2023]:
@@
Appendix C Adaptation Evaluation and Model details
@@
Falcon Instruct: A version of Falcon 7B finetuned on the Baize [Xu et al., 2023], GPT4All [Anand et al., 2023], GPTeacher [Teknium1, 2023], and Refined-Web English [Penedo et al., 2023] datasets. Retrieved from https://huggingface.co/tiiuae/falcon-7b-instruct.
RPJ-INCITE Chat: A version of RPJ-INCITE 7B finetuned on the OASST1 [Köpf et al., 2023] and Dolly V2 [Conover et al., 2023] datasets. Retrieved from https://huggingface.co/togethercomputer/RedPajama-INCITE-7B-Chat.
Llama-2 Chat: A version of Llama 2 7B finetuned on a mixture of instruction datasets and further trained with RLHF. We refer the reader to Touvron et al. [2023b] for further details.
Tülu 2: A version of Llama 2 7B finetuned on a mixture of instruction datasets (the Tülu 2 mix). We refer the reader to Ivison et al. [2023] for further details.
@@
OLMo +SFT+DPO: OLMo +SFT further trained with DPO on the UltraFeedback dataset [Cui et al., 2023].
@@
MMLU: We use the official MMLU [Hendrycks et al., 2021] evaluation script and prompts available at https://github.com/hendrycks/test, with modifications to allow for batch processing. We evaluate using 0 few-shot examples, following the original setup of MMLU. We report average accuracy across test examples.
ToxiGen: We follow the setup in Touvron et al. [2023b], but use the original set of prompts from Hartvigsen et al. [2022], which are designed to elicit toxic generations for certain groups. We take only the prompts designed to produce toxic language (‘hateful’ prompts) and use 500 prompts per group to reduce evaluation costs. For base language models, we pass in the original ToxiGen prompts unchanged and greedily decode up to the first new line (or a maximum of 512 tokens). For instruction-tuned models, we place the prompt in the corresponding template, and ask the model to complete the prompt, until the model generates a stop token (or a maximum of 512 tokens). We pass the generated text into a roberta-large model trained to detect toxic content finetuned as part of Hartvigsen et al. [2022]282828https://huggingface.co/tomh/toxigen_roberta. We then report the percentage of generations deemed toxic by the classifier.
TruthfulQA: Following Touvron et al. [2023b], we mainly use the generation setting of TruthfulQA [Lin et al., 2022]. The TruthfulQA dataset contains 818 questions, which are used to prompt the tested model to generate answers. We use the default QA prompt format with 6 in-context QA examples. We follow the official script in their official implemention292929https://github.com/sylinrl/TruthfulQA/ to do greedy decoding and answer postprocessing. We train two LLaMA 2-based classifiers for judging the truthfulness and informativeness of the model response, due to the deprecation of GPT-3 making exact replication of the original TruthfulQA evaluation infeasible. We find that the LLaMA 2 judges are generally able to match the performance of the original GPT-3-based judges used by Lin et al. [2022]. We report the rate of the responses being truthful and informative (% Informative and Truthful) following Touvron et al. [2023b]. We only report the % Informative and Truthful as our primary metric.
AlpacaEval: We use the package provided by Li et al. [2023], following the default setup which asks the evaluated model to generate responses for 805 prompts and employ GPT-4 to compare the response with Davinci-003. We employ the “alpaca_eval_gpt4” annotator. We allow the evaluated model to generate up to 2048 tokens, without specifying special stop sequences. The reported win-rate is the percentage of model generations that GPT-4 reports as being preferred over the generations from Davinci-003.

### v4

@@ Appendix D Adaptation Training Details
Data: Tülu V2 SFT mix, resplit such that long conversations are split into 2048-token chunks and replacing the hardcoded split with data about OLMo. Data is publically available.151515https://huggingface.co/datasets/allenai/tulu-v2-sft-mixture-olmo-2048
After instruction finetuning, we then use the following hyperparameters for DPO training, following Ivison et al. (2023):
@@
Data: A modified form of UltraFeedback (Cui et al., 2023), with TruthfulQA prompts removed. We used the ‘fixed’ variant released by Argilla, which uses the average of GPT-generated aspect-based scores to determine chosen and rejected pairs.161616https://huggingface.co/datasets/argilla/ultrafeedback-binarized-preferences-cleaned
Appendix E Adaptation Evaluation and Model details
@@
Falcon Instruct: A version of Falcon 7B finetuned on the Baize (Xu et al., 2023), GPT4All (Anand et al., 2023), GPTeacher (Teknium1, 2023), and Refined-Web English (Penedo et al., 2023) datasets. Retrieved from https://huggingface.co/tiiuae/falcon-7b-instruct.
RPJ-INCITE Chat: A version of RPJ-INCITE 7B finetuned on the OASST1 (Köpf et al., 2023) and Dolly V2 (Conover et al., 2023) datasets. Retrieved from https://huggingface.co/togethercomputer/RedPajama-INCITE-7B-Chat.
Llama-2 Chat: A version of Llama 2 7B finetuned on a mixture of instruction datasets and further trained with RLHF. We refer the reader to Touvron et al. (2023b) for further details.
Tülu 2: A version of Llama 2 7B finetuned on a mixture of instruction datasets (the Tülu 2 mix). We refer the reader to Ivison et al. (2023) for further details.
@@
OLMo+SFT+DPO: OLMo+SFT further trained with DPO on the UltraFeedback dataset (Cui et al., 2023).
@@
MMLU: We use the official MMLU (Hendrycks et al., 2021) evaluation script and prompts available at https://github.com/hendrycks/test, with modifications to allow for batch processing. We evaluate using 0 few-shot examples, following the original setup of MMLU. We report average accuracy across test examples.
ToxiGen: We follow the setup in Touvron et al. (2023b), but use the original set of prompts from Hartvigsen et al. (2022), which are designed to elicit toxic generations for certain groups. We take only the prompts designed to produce toxic language (‘hateful’ prompts) and use 500 prompts per group to reduce evaluation costs. For base language models, we pass in the original ToxiGen prompts unchanged and greedily decode up to the first new line (or a maximum of 512 tokens). For instruction-tuned models, we place the prompt in the corresponding template, and ask the model to complete the prompt, until the model generates a stop token (or a maximum of 512 tokens). We pass the generated text into a roberta-large model trained to detect toxic content finetuned as part of Hartvigsen et al. (2022).171717https://huggingface.co/tomh/toxigen_roberta We then report the percentage of generations deemed toxic by the classifier.
TruthfulQA: Following Touvron et al. (2023b), we mainly use the generation setting of TruthfulQA (Lin et al., 2022). The TruthfulQA dataset contains 818 questions, which are used to prompt the tested model to generate answers. We use the default QA prompt format with 6 in-context QA examples. We follow the official script in their official implemention181818https://github.com/sylinrl/TruthfulQA/ to do greedy decoding and answer postprocessing. We train two LLaMA 2-based classifiers for judging the truthfulness and informativeness of the model response, due to the deprecation of GPT-3 making exact replication of the original TruthfulQA evaluation infeasible. We find that the LLaMA 2 judges are generally able to match the performance of the original GPT-3-based judges used by Lin et al. (2022). We report the rate of the responses being truthful and informative (% Informative and Truthful) following Touvron et al. (2023b). We only report the % Informative and Truthful as our primary metric.
AlpacaEval: We use the package provided by Li et al. (2023), following the default setup which asks the evaluated model to generate responses for 805 prompts and employ GPT-4 to compare the response with Davinci-003. We employ the “alpaca_eval_gpt4” annotator. We allow the evaluated model to generate up to 2048 tokens, without specifying special stop sequences. The reported win-rate is the percentage of model generations that GPT-4 reports as being preferred over the generations from Davinci-003.